# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

### v2

@@ Abstract
Grammatical error correction (GEC) aims to improve quality and readability of texts through accurate correction of linguistic mistakes. Previous work has focused on high-resource languages, while low-resource languages lack robust tools. However, low-resource languages often face problems such as: non-standard orthography, limited annotated corpora, and diverse dialects, which slows down the development of GEC tools. We present a study on GEC for Zarma, spoken by over five million in West Africa. We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs). We evaluated them using a dataset of more than 250,000 examples, including synthetic and human-annotated data. Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95. 82% and a suggestion accuracy of 78. 90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections. The rule-based method was effective for spelling errors but failed on complex context-level errors. LLMs—MT5-small—showed moderate performance. Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
@@
Mamadou K. KEITA1, Christopher Homan3, Marcos Zampieri4, Adwoa Bremang2, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1PAAIR Institute, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
@@ 1 Introduction
Grammatical error correction (GEC) is important to improving text quality in different scenarios. High-resource languages have well-developed GEC tools (Napoles et al., 2019), but low-resource languages have few resources. Zarma, a Nilo-Saharan language spoken widely in West Africa (Lewis et al., 2016), faces these challenges.
Zarma texts lack standardized orthography and large annotated datasets. Developers need approaches that handle minimal resources and non-standard writing. Large language models (LLMs) hold promise—as they can handle various languages and tasks (Brown et al., 2020)—but performance on low-resource languages can be limited. They rely heavily on data, which is scarce for Zarma.
We investigate three GEC approaches for Zarma: a rule-based method, a machine translation (MT)-based model, and LLMs. We also test with Bambara, another West African language, to confirm generalization. This paper addresses two key questions:
@@
  • A dataset of more than 250,000 synthetic and human-annotated Zarma examples.
@@
  • Planned release of trained models to promote future research.
@@ 2 Related Work
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. The complexity of balancing precision and recall on synthetic data remains an open problem.
3 Proposed Methods
@@ 3.1 GEC with LLMs
LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for using LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
@@ 3.1.1 Implementation
Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
@@
We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model’s ability to align corrupted text with corrected forms.
@@ 3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. We adapt M2M100 for GEC by training it to translate corrupted Zarma text into corrected versions. This approach uses a synthetic corpus generated with a noise script (Section 4). The script introduces various errors, ensuring that the training data effectively represent realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
@@ 3.3 Rule-based GEC for Zarma
The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions. This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma’s linguistic intricacies.
@@
The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format. This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma’s grammatical essence despite its absence of a writing standard.
@@
When an error is detected, either because the Bloom Filter returned false or the grammatical check identified an issue, the Correction Algorithm is activated. This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word. For illustration, given a corrupted sentence "A sindq biri," the Levenshtein algorithm’s operation can be represented as:
Corrupted Sentence: "A sindq $\begin{array}[]{c}\text{{Corrupted Sentence}: "A sindq biri"}\\ \downarrow\\ \text{{LevSuggest}}("sindq")\\ \downarrow\\ \{\text{"sind"},\text{"sinda"}\}\end{array}$ (3)
Here, "sindq" undergoes comparison with the lexicon, and "sind" and "sinda" are suggested based on their proximity in terms of minimal edit distance. In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies. This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
@@ 4 Data
We created a combined dataset of over 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ”gold” dataset. We split it into 80% training, 10% validation, and 10% testing.
4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, We applied a noise script to the Feriji dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, to mimick plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
We applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
$CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$.
For example, consider the original sentence SZ = A go koy fuo” (He is going home”). Applying a substitution operation $\delta$ that changes ’go’ to ’ga’ yields the corrupted sentence CSZ = ’A ga koy fuo’. While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
$\textbf{SZ}="A\,go\,koy\,fuo"(I\ am\ going\ home)$ (5)
$\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
$\textbf{CSZ}="A\,ga\,koy\,fuo"(I\ will\ go\ home)$
[…]
Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset. A snapshot of the synthetic data is provided in Table 1.
@@ 4.1 Human-Annotated Data
The second dataset was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the human-annotated dataset is shown in Table 2
@@
Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
@@ 6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings such as Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
@@ 6.1 Automatic Evaluation Results
In the AE, we employed two sets of metrics to evaluate our techniques. The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012). However, these metrics are not applicable to the rule-based method due to its deterministic nature and the way these metrics are designed, which favor probabilistic models that produce varying outputs. The second set focused on spelling correction metrics—measuring the detection rate and suggestion accuracy. The results are presented in Table 3.
@@
GLEU Detection Correction
Rule-based - - 100% 96.27%
Gemma 2b 0.4572 0.528 76.19% 43.28%
MT5-small 0.6203 0.731 90.62% 57.15%
M2M100 0.7896 0.914 95.82% 78.90%
@@ 6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
@@ 6.1.2 Spelling Correction Metrics
For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4). We measured two key metrics:
@@
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This high performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.

### v3

@@ Abstract
Grammatical error correction (GEC) aims to improve text quality and readability. Previous work on the task focused primarily on high-resource languages, while low-resource languages lack robust tools. To address this shortcoming, we present a study on GEC for Zarma, a language spoken by over five million people in West Africa. We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs). We evaluated GEC models using a dataset of more than 250,000 examples, including synthetic and human-annotated data. Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections. The rule-based method was effective for spelling errors but failed on complex context-level errors. LLMs—Gemma 2b and MT5-small—showed moderate performance. Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
@@
Mamadou K. Keita1, Adwoa Bremang2, Huy Le1 Dennis Owusu2, Marcos Zampieri3, Christopher Homan1 1Rochester Institute of Technology, 2Ashesi University, 3George Mason University
@@ 1 Introduction
Grammatical error correction (GEC) is important to improving text quality in different scenarios. High-resource languages have well-developed GEC tools (Bryant et al., 2023; Ng et al., 2025) while low-resource languages have few resources. Zarma—a Nilo-Saharan language spoken by over five million people, primarily in Niger, where it is a major national language, as well as in Nigeria, Mali, and Burkina Faso (Lewis et al., 2016)—faces these challenges.
Zarma lacks a standardized orthography and large annotated datasets. To create language technology for low-resource languages like Zara, developers need approaches that handle minimal resources and non-standard writing Tapo et al. (2020, 2025). Large language models (LLMs) hold promise as they are capable of handling various languages and tasks (Brown et al., 2020), but performance on very low-resource languages can be limited. LLMs rely heavily on data, which is scarce for Zarma.
To address these important challenges in low-resource NLP, we investigate GEC approaches for Zarma. We evaluate three approaches: a rule-based method, a machine translation (MT)-based model, and LLMs. We also test with Bambara, another West African language, to confirm generalization. This paper addresses three key research questions (RQs):
@@
  2. 2.
    A dataset of 250,000 synthetic and human-annotated Zarma examples  [^1] The latest version of the dataset can be found here: 27Group/noisy_zarma.
  3. 3.
@@ 2 Related Work
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the No Error Left Behind approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a translation task, where the source language is the incorrect sentence and the target language is the corrected sentence. The complexity of balancing precision and recall on synthetic data remains an open problem.
3 Methods
@@ 3.1 GEC with LLMs
LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel and others, 2020). This section outlines our methodology for using LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models' few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
@@ 3.1.1 Implementation
[…]
This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei and others, 2021).
@@
We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model's ability to align corrupted text with corrected forms.
@@ 3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. We adapt M2M100 for GEC by training it to translate corrupted Zarma text into corrected versions. This approach uses a synthetic corpus generated with a noise script (Section 4). The script introduces various errors, ensuring that the training data effectively represent realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 8, 5, and 6.
@@ 3.3 Rule-based GEC for Zarma
The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions. This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma's linguistic intricacies.
@@
The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format. Zarma is classified as an agglutinative language, meaning it depends heavily on attaching suffixes to root words to indicate grammatical features like plurality and definiteness (e.g., the definite plural suffix -ey). This morphology imposes specific constraints on our rule-based approach: the rules must be robust enough to parse and validate these suffix chains rather than treating words as static units. This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma's grammatical essence despite its absence of a writing standard.
@@
When an error is detected, either because the Bloom filter returned false or the grammatical check identified an issue, the Correction Algorithm is activated. This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word. For illustration, given a corrupted sentence ``A sindq biri,'' the Levenshtein algorithm's operation can be represented as:
$\begin{array}[]{c}\text{{Corrupted Sentence}: ``A sindq biri''}\\ \downarrow\\ \text{{LevSuggest}}(``sindq^{\prime\prime})\\ \downarrow\\ \{\text{``sind''},\text{``sinda''}\}\end{array}$ (3)
Here, ``sindq'' undergoes comparison with the lexicon, and ``sind'' and ``sinda'' are suggested based on their proximity in terms of minimal edit distance. In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies. This mechanism ensures the algorithm's efficiency in providing appropriate corrections, thus enhancing the GEC tool's overall accuracy for Zarma text.
@@ 4 Data
To address the scarcity of annotated data in Zarma, we created a combined dataset of 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ``gold'' dataset. The synthetic dataset contains 248,000 sentence pairs; the human-annotated dataset includes 2,000 sentence pairs. We split this combined dataset into 80% training, 10% validation, and 10% testing sets.
4.1 Synthetic Data
@@
We applied a noise script to the Feriji dataset (Keita et al., 2024) to generate synthetic errors. This script introduces typographical and grammatical errors into Zarma sentences—errors that mimic human mistakes. To ensure these errors reflect real-world patterns, we worked with five native Zarma speakers—the same individuals who later conducted the manual evaluation (ME) as described in Section B. We provided them with audio recordings of 46 Zarma sentences from everyday conversations, which they transcribed into written Zarma—with instructions to not attempt to correct any mistake. All typed in French at an average of 38 words per minute using an AZERTY keyboard. We compared their transcriptions with the correct Zarma text and analyzed the frequency of error types. The study revealed common errors: insertions (e.g., ``Irikoy'' to ``Irikoyu''), and deletion (e.g., ``alfarey'' to ``alfary''); with frequencies of 23% and 12%, respectively. Minor transpositions and double vowels were noticed too—less than 2% and 7% respectively. The AZERTY keyboard layout and typing speed influence error types such as insertions due to adjacent key proximity—e.g, `u' near `y' in `Irikoyu'). Also, the average error per sentence was 1 error for short sentences (1-5 words), 3 errors for long sentences (5+ words).
The noise function $\mathbf{N}$ consists of four operations: deletions $\mathbf{\delta}$, insertions $\mathbf{\mu}$, substitutions $\mathbf{\sigma}$, and transpositions $\mathbf{\tau}$. For example, applying a substitution operation to the original sentence $\mathbf{SZ}=\textit{A go koy fuo}$ (``He is going home'') might yield $\mathbf{CSZ}=\textit{A ga koy fuo}$ (``He will go home''), introducing a grammatical error by changing the tense.
$\mathbf{SZ}=``A\,go\,koy\,fuo^{\prime\prime}$ (4)
$\mathbf{N}(\mathbf{SZ})=\mathbf{SZ}\xrightarrow{\sigma(``go^{\prime\prime}\to``ga^{\prime\prime})}``A\,ga\,koy\,fuo^{\prime\prime}$
$\mathbf{CSZ}=``A\,ga\,koy\,fuo^{\prime\prime}$
$\mathbf{SZ}\xrightarrow{N}\mathbf{Correct}\,\mathbf{CSZ}$
Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset. A snapshot of the synthetic data appears in Table 1.
@@ 4.2 Human-Annotated Data
The human-annotated dataset was curated by native Zarma speakers who introduced logical and grammatical (LG) errors into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly logical errors—and their corresponding corrections, enhancing the models' capacity to generalize and accurately correct unseen texts in Zarma. An example from the human-annotated dataset is shown in Table 2
@@
Souba, Ay koy Niamey Souba, Ay ga koy Niamey ``Souba'' means tomorrow, and therefore the tense must be in the future using the future tense marker ``ga'' after the subject ``Ay.''
@@ 6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section A), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and ME by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings such as Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
@@ 6.1 Automatic Evaluation Results
In the AE, we employed two sets of metrics to evaluate our techniques. The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012). The second set focused on spelling correction metrics—measuring the detection rate, suggestion accuracy, and false positive rate. The results are presented in Table 3.
@@
GLEU Detection Correction False Positive
Rule-based 0.3124 0.401 100% 96.27% 2.5%
Gemma 2b 0.4572 0.528 76.19% 43.28% 15.3%
MT5-small 0.6203 0.731 90.62% 57.15% 8.7%
M2M100 0.7896 0.914 95.82% 78.90% 4.2%
@@ 6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models' performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The rule-based method achieved modest scores (GLEU: 0.3124, M2: 0.401), reflecting its effectiveness in spelling correction but limitations in handling complex grammatical errors. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its superior performance among the tested models. The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
@@ 6.1.2 Spelling Correction Metrics
For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4). We measured three key metrics:
@@
  • False Positive Rate [^2] While we did not evaluate the models on entirely correct sentences, our False Positive Rate reflects over-corrections within mixed correct-and-incorrect datasets, approximating real-world GEC scenarios where fully pristine text is rare.: The percentage of correct words incorrectly flagged as errors. Here, a false positive is counted not only when an entire sentence is flagged, but also when parts of a sentence that are correct are unnecessarily altered. In other words, if a sentence contains both correct and incorrect segments, any modification of the correct segments is considered a false positive.
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%, with a low false positive rate of 2.5%. This high performance is primarily due to its dictionary—built from the Feriji Dataset—and predefined rules, which allow it to effectively flag and correct errors within its scope while rarely misidentifying correct words. The M2M100 model also performed well, with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and a false positive rate of 4.2%, indicating its capability to identify and correct spelling errors in a with minimal over-correction. The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies, coupled with higher false positive rates. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28% and a false positive rate of 15.3%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15% and a false positive rate of 8.7%. These results suggest that while LLMs can correct some spelling errors, they are less effective than the MT-based and rule-based methods in this task and are more prone to incorrectly flagging correct words. The reasons of this poor performance of LLMs is due to the fact they were not pre-trained on Zarma, thereby limiting their accuracies in Zarma related tasks.
6.2 Performance on Synthetic and Human-Annotated Subsets
To further explore how each method handles different data sources, we tested performance on the synthetic and human-annotated subsets described in Section 4. Table 4 presents the results side by side.
Methods Synthetic Subset Human-Annotated Subset
GLEU GLEU
Rule-based 0.3150 0.405 0.2900 0.375
Gemma 2b 0.4601 0.531 0.4213 0.492
MT5-small 0.6257 0.738 0.5812 0.692
M2M100 0.7923 0.917 0.7618 0.891
Table 4: Performance metrics on synthetic and human-annotated subsets
Table 4 reveals a pattern—all models, including the rule-based method, score lower on the human-annotated subset than on the synthetic subset. For the rule-based method, GLEU drops from 0.31 to 0.2900 and M² from 0.40 to 0.375, indicating its reduced capability with complex logical errors present in human-annotated data. This difference arises because human-annotated data includes complex logical and context-level errors—challenges beyond the simpler typographical errors in the synthetic set. Nevertheless, M2M100 maintains its lead across both subsets. Its GLEU drops from 0.7923 to 0.7618 and M² from 0.917 to 0.891, yet these scores still outperform Gemma 2b and MT5-small by 5%–10%. This consistency highlights M2M100’s strength in tackling diverse error types.